OpenStudy (anonymous):

How do I solve x²-i=0 using the n(th) roots of a complex number in trigonometry?

OpenStudy (anonymous):

\[\sqrt[n]{z}=\sqrt[n]{|z|}(cosArg(z)/n+i \sin Arg(z)/n)\]

OpenStudy (anonymous):

is x a factor or can I just ignore x then?

OpenStudy (anonymous):

x is equal sqrt (i)

OpenStudy (anonymous):

so I have standard form sqrt(i)? I dont need to do anything with x?

OpenStudy (anonymous):

write it in this way:\[x ^{2}=i \rightarrow x =\sqrt{i}\] so just use the formula to find sqrt of i

OpenStudy (anonymous):

x is what you trying to find....

OpenStudy (anonymous):

ok ok I think I got it, the x was throwing me off thanks man

OpenStudy (anonymous):

I'm not used to the x being in the equation thanks a lot man

OpenStudy (anonymous):

you wellcome

OpenStudy (experimentx):

x²=i = e^i pi/2 x = +-e^ i pi/4

OpenStudy (anonymous):

thanks